# Standardizing the Measurement of Text Diversity: A Tool and a Comparative Analysis of Scores

### v1

@@ Abstract
The diversity across outputs generated by large language models shapes the perception of their quality and utility. Prompt leaks, templated answer structure, and canned responses across different interactions are readily noticed by people, but there is no standard score to measure this aspect of model behavior. In this work we empirically investigate diversity scores on English texts. We find that computationally efficient compression algorithms capture information similar to what is measured by slow to compute $n$-gram overlap homogeneity scores. Further, a combination of measures—compression ratios, self-repetition of long $n$-grams and Self-BLEU and BERTScore—are sufficient to report, as they have low mutual correlation with each other. The applicability of scores extends beyond analysis of generative models; for example, we highlight applications on instruction-tuning datasets and human-produced texts. We release a diversity score package to facilitate research and invite consistency across reports.
@@ 1 Introduction
Evaluation of LLM-generated texts is typically done with respect to accuracy or factuality, e.g., as measured via entailment (Tang et al., 2023), or text quality aspects such as coherence and fluency (e.g., estimated using LLMs as evaluators as in Liu et al. 2023). For summarization tasks where references are available, the similarity of generated outputs to these is also often measured. A complementary dimension of model performance is diversity, i.e., how much “boilerplate” content is repeated across LLM outputs.
The diversity of generated outputs is intuitively important: A model prone to repeating specific sentence constructions or boilerplate turns of phrase across its outputs will likely be deemed lower quality than an LLM capable of more diverse generations, all else being equal. Table 1 shows example repetitions over a news summarization dataset. We provide detailed examples including which model produced each set of examples in Appendix A.1. In this work we first analyze commonly reported diversity scores over English language outputs from several LLMs, and identify a few practical, (mostly) independent scores that characterize repetition. We also release diversity, an open-source Python package that can be used to explore and evaluate the diversity of a generated text datasets. [^1] https://pypi.org/project/diversity/
Model Token Repetition Text Pattern-Matched Text
StableLM ”The article also notes that…” 41/500 DT NN VBZ DT JJ NN 84/500
”The article also mentions that…” 17/500 ”The article discusses the recent debate…”
”The article notes that the…” 15/500 ”The book provides a helpful guide…”
”The article concludes by stating…” 14/500 ”The article discusses the controversial penalty…”
”The author also notes that…” 11/500 ”The article discusses the illicit market…”
FlanT5 ”The boy, who cannot be…” 5/500 NNP NNP DT JJ NN VBD 37/500
”The study, published in the…” 7/500 ”Christopher Barry, a 53-year-old man, was…”
”The body of a man who…” 3/500 ”Charles Collins, a 28-year-old man, saved…”
”A man has been charged with…” 5/500 ”Damian Parks, a 22-year-old student, went…”
”… was last seen at a…” 4/500 ”Lynn Fast, a 21-year-old mother, claimed…”
Llama-2 ”The article discusses the…” 15/500 NNP NN NNP NNP VBZ VBN 42/500
”The article also mentions that…” 28/500 ”American conductor Marin Alsop has become…”
”The article concludes by noting…” 6/500 ”England striker Andy Cole has warned…”
”The article is about the…” 19/500 ”Barcelona defender Dani Alves has announced…”
”According to the article, a…” 16/500 ”England economist Neil Haldane has said….”
@@
Table 1: Examples of exact text-match and repeating part-of-speech patterns in 500 model generated summaries of news articles from the CNN/DM dataset. The number of times the pattern occurs is shown in parenthesis. These patterns appear at a higher frequency in the generated outputs than in the original input data. Different models are characterized by a different set of repeated patterns.
2 Background
We investigate automatic approaches for characterizing the diversity of outputs from LLMs. Our work is motivated by qualitative observations of such outputs, in which it is easy to notice undesirable repetition of formulaic responses.
Lack of diversity may result from repetition of long stretches of text or be due to subtle distributional patterns (Holtzman et al., 2019; Meister et al., 2022; 2023a). We focus on scores that are likely to capture overt repetition across outputs, and leave for future work similar analysis of semantic and structural diversity scores (Bär et al., 2012).
Image captioning and dialog are text generation tasks with a robust body of work quantifying the diversity of produced texts. In both domains, prior work has documented that models tend to produce the same text for different contexts. Li et al. (2016) find that four phrases account for about a third of all turns produced by a conversational agent. Devlin et al. (2015) report that more than half of the automatic captions are repeated verbatim for different images and most of these captions are seen as exact strings in training. Self-repetition (Salkar et al., 2022), i.e., an exact repetition of the same $n$-gram longer than four across different summaries is a way to adapt for a similar analysis in tasks where the generated text is longer, so exact matches rarely occur but repetition is common, especially relative to the training data (Wang et al., 2023a).
@@ 2 Background
The above observations speak to the need for a standardized, easy to use method to quantify diversity. In this work we propose such standardization and show that compression ratio is a fast, convenient to compute score which is sufficient to capture the information in all token/type ratio related alternatives. However, we also find that compression ratios—and all scores considered—are moderately to strongly correlated with the length of texts, complicating interpretation. Many comparisons remain meaningful when accompanied with information about length, but in others no conclusions can be reliably drawn.
The association between length and measures of diversity is well-established in corpus linguistics (see detailed discussion in Brysbaert et al. 2016). The number of unique words in a corpus is a power function of the total words seen, where the power is less than 1. The number of total words grows linearly, while the number of unique words is sublinear, so longer texts have more repetitions of unique words and $n$-grams than shorter texts (Covington & McFall, 2010; McCarthy & Jarvis, 2010).
LLM generations can lead to reduced text diversity in a few ways. Guo et al. (2023) study the effect of consecutive rounds of distillation, in which a language model produces the data on which the next language model is trained. They report dramatic reduction in diversity over 10 iterations. However, this study did not report the length of text produced in consecutive distillation rounds. Given our findings, it would be prudent to check if output lengths remain comparable (or shorter). Padmakumar & He (2023a) show that when people write with the aid of an LLM (e.g., instructGPT) they produce less diverse writing than when they do not. Here we ascertain that these results are independent of length, indicating that there is a genuine reduction in diversity (rather than merely affecting lengths which in turn influence diversity measures).
Work across use-cases has shown that reducing the repetition—or, equivalently, increasing the diversity—in training data yields higher quality models. De-duplicating pre-training data leads to more efficient pre-training and better models that do not repeat pre-training data directly (Lee et al., 2022; Abbas et al., 2023). Removing fine-tuning summaries with repeated content improves summarization performance (Choubey et al., 2023). Given this, instruction diversity used for instruction tuning LLMs will likely also have implications for mode performance. We assess the diversity of instructions in common instruction tuning datasets.
3 A Smorgasbord of Text Diversity Scores
@@
The variety of scores used to measure diversity across a corpus of texts derive from two core ideas: Computing average similarity between pairs of outputs produced by the same model for different inputs, and computing variants of token/type ratio. The former are adapted from common approaches to text generation evaluation by comparing with references, using standard measures of pairwise similarity; the latter track the diversity of vocabulary measured as the ratio of unique words to total words produced, with the outputs from a model concatenated into a single text.
We first describe each score, and then present insights regarding their mutual redundancy. We also consider their required run-times, which are lengthy for some metrics and may render them impractical for analysis of a large number of outputs. All scores are defined for a set of generated texts $D$, each conditioned on its respective input.
@@
The quality of text in machine translation, summarization, and image captioning is often reported in terms of overlap with a reference text. This idea can be adapted to measure diversity across different outputs by using one generated text as a reference and measuring the similarity of other outputs against this. Self-BLEU measures similarity between all text pairs in $D$ using BLEU as the similarity score (Zhu et al., 2018). BLEU can be replaced with an arbitrary similarity score, e.g., ROUGE or BERTScore. These variants are called homogenization scores and have recently been used to compare the diversity of texts produced under several conditions (Padmakumar & He, 2023a).
@@
Here the similarity score of choice is ROUGE-L (Lin & Och 2004; Eq. 1). This quantifies overlap in terms of longest common sub-sequences between all pairs of text in a corpus instead of the fixed $n$-gram size used in other ROUGE variants:
@@
Self-repetition was introduced to measure the tendency of LMs to repeat long $n$-grams across different outputs (Salkar et al., 2022).
@@
The token-type ratio for a text is the unique token count divided by the total count of tokens. Moving Average Token Type Ratios (MATTRs) was introduced as a way to measure the lexical dynamics across a text which is insensitive to text length. The score captures the repetition of a given word in segments of text and does not explicitly account for longer repeated sequences (Covington & McFall, 2010).
@@
NGD extends the idea of token-type ratio to longer $n$-grams (Padmakumar & He, 2023a; Meister et al., 2023b; Li et al., 2023). It is defined as a ratio of the unique $n$-gram counts to all $n$-gram counts:
@@
The probability of text under a Hypergeometric Distribution D (HD-D) is an another measure of lexical diversity (McCarthy & Jarvis, 2010). [^2] For both HD-D and MATTR, we use the implementation provided in the lexical-diversity package (https://pypi.org/project/lexical-diversity/). HD-D does not capture repetition of sub-sequences.
@@
The diversity scores introduced so far are all a function of the number of repeated substrings across outputs. Some measure these over pairs of generated texts, others are computed for a concatenation of all outputs into a single text. Text compression algorithms are designed to identify redundancy of sequences of variable length in text.
We use gZip to compress the concatenated text of all outputs generated by a model. The compression ratio is then the ratio between the size of the compressed file to that of the original file. High compression ratios imply more redundancy:
@@
4 Data and Models
@@
We compute diversity scores for the outputs of six instruction tuned models on the CNN/DailyMail (Hermann et al., 2015) and XSUM (Narayan et al., 2018) English news summarization datasets. The models are: Llama-2 (Touvron et al., 2023a), GPT-4 (OpenAI, 2023), FlanT5-XXL (Longpre et al., 2023), StableLM (Taori et al., 2023; Chiang et al., 2023; Anand et al., 2023), Mistral (Jiang et al., 2023), and StableBeluga (Touvron et al., 2023b; Mukherjee et al., 2023). [^4] All models—except GPT-4—downloaded from HuggingFace (https://huggingface.co/models). We selected these models to cover a range of availability (open- and closed-source), and architectures (encoder-decoder, decoder-only). [^5] We use prompts for summarization provided by each model, where available. See Appendix A.2. We also measure the diversity of the input news articles, the first three sentences of the articles, and the reference summaries. The lengths of texts vary considerably by source, for reference and model-produced text alike, so we also note average lengths when reporting diversity.
5 Text Length as Confounder
To keep computational time and costs manageable, we randomly sample 500 inputs from CNN/DailyMail and XSUM for analysis. Table 3 reports diversity scores for the outputs generated by the six zero-shot LLMs for these inputs.
CR
CR: POS
NGD
Self-
Rep.
Hom.
(R-L)
Hom.
Self-
BLEU
MATTR HD-D
0.83 0.695 0.885 0.87 0.841 0.921 0.991 0.799 0.654

### v2

# Standardizing the Measurement of Text Diversity: A Tool and Comparative Analysis

@@ Abstract
The diversity across outputs generated by LLMs shapes perception of their quality and utility. High lexical diversity is often desirable, but there is no standard method to measure this property. Templated answer structures and “canned” responses across different documents are readily noticeable, but difficult to visualize across large corpora. This work aims to standardize measurement of text diversity. Specifically, we empirically investigate the convergent validity of existing scores across English texts, and we release diversity, an open-source Python package [^1] https://pypi.org/project/diversity/ for measuring and extracting repetition in text. We also build a platform [^2] https://ai-templates.app based on diversity for users to interactively explore repetition in text. We find that fast compression algorithms capture information similar to what is measured by slow-to-compute $n$-gram overlap homogeneity scores. Further, a combination of measures—compression ratios, self-repetition of long $n$-grams, and Self-BLEU—are sufficient to report, as they have low mutual correlation with each other.
[…]
Standardizing the Measurement of Text Diversity:
A Tool and Comparative Analysis
Chantal Shaib1thanks: Partial work completed while at Adobe Research.    Joe Barrow3footnotemark:     Jiuding Sun1    Alexa F. Siu2 Byron C. Wallace1  Ani Nenkova2 1Northeastern University, 2Adobe Research, 3Pattern Data shaib.c@northeastern.edu
@@ 1 Introduction
Figure 1: Part-of-speech pattern visualization with the diversity app. Users upload a corpus, and the UI returns frequent part-of-speech sequences. Users can navigate through different pattern lengths of $n=(2,10)$.
Figure 2: (A) Users start by uploading their own dataset on the right or by selecting one of the existing demo datasets on the left. Once uploaded, users can (B) interactively visualize part-of-speech patterns in the data, (C) interactively search for exact repeated text matches, and (D) calculate lexical diversity metrics.
@@
Evaluation of LLM-generated texts is typically done with respect to accuracy or factuality, e.g., as measured via entailment (Tang et al., 2023), or text quality aspects such as coherence and fluency (e.g., estimated using LLMs as evaluators as in Liu et al. 2023). For tasks where reference summaries are available, the similarity of generated outputs to these is also often measured (e.g., using ROUGE scores; Lin and Och, 2004). A complementary dimension of model performance is diversity, or how much “boilerplate” content is repeated across LLM outputs. For example, prompted to summarize news articles, Llama-2 (Touvron et al., 2023b) tends to generate text that begins with “The article discusses that…", “The article notes that…". Table 1 illustrates example repetitions from various models over a news summarization dataset.
Token Repetition Text
"The article also notes that…" (41/500 )
"The article also mentions that…" (17/500)
"The article notes that the…" (15/500)
"The article discusses the…" (15/500)
"The article also mentions that…" (28/500)
"The article is about the…" (19/500)
"According to the article, a…" (16/500)
Pattern-Matched Text
DT NN VBZ DT JJ NN (84/500)
"The article discusses the recent debate…"
"The book provides a helpful guide…"
"The article discusses the controversial penalty…"
"The article discusses the illicit market…"
NNP NNP DT JJ NN VBD (37/500)
"Christopher Barry, a 53-year-old man, was…"
"Charles Collins, a 28-year-old man, saved…"
"Damian Parks, a 22-year-old student, went…"
"Lynn Fast, a 21-year-old mother, claimed…"
@@
Table 1: Exact text match and repeated part-of-speech patterns across 500 LLM-generated summaries of news articles from CNN/DM. The number of times the pattern occurs is reported parenthetically.
There is a distinct lack of standardization in reporting diversity in machine learning datasets (Zhao et al., 2024). Evaluating diversity is complicated by a lack of tools enabling quantification and exploration of repetitive text in large corpora. This work primarily aims to standardize such reporting, focusing on text-based corpora. We introduce diversity, an open-source Python package for evaluating diversity.1 In addition to this package, we provide a standalone web-based tool that allows users to visualize repetition in their corpus.2 This UI provides an intuitive, efficient tool for text analysis, allowing users to: (i) View repetitive text and part-of-speech $n$-grams of varying lengths; (ii) Quickly compute diversity metrics, and; (iii) Interactively highlight and match repetition in documents. To facilitate future exploration and measurement of lexical diversity, both package [^3] https://github.com/cshaib/diversity and UI code [^4] https://github.com/cshaib/diversity_demo are available under the Apache 2.0 license.
To validate existing diversity metrics, we analyze them over English language outputs from several LLMs. This identifies a few practical, (mostly) independent scores that characterize repetition. We emphasize text length as an important confounder when assessing diversity, and examine diversity in downstream datasets such as instruction tuning. When measuring lexical diversity, many comparisons remain meaningful when accompanied with information about length, but absent this variable no reliable conclusions can be drawn.
@@
In addition to assessing the information overlap of existing metrics, we show that compression ratio—a measure of document compression relative to original size—is a fast, easy to compute score that is sufficient to capture the information in all token/type ratio related alternatives. However, we also find that compression ratios (and all scores considered) are moderately to strongly correlated with text length, complicating interpretation.
Our contributions are as follows. (1) We introduce diversity, a Python package implementing diversity metrics. (2) We host and release source code to a user interface to explore repetition and diversity. (3) We evaluate the convergent validity of existing lexical diversity metrics and highlight compression ratios as efficient measures of diversity. (4) We demonstrate use cases of our package in NLP research via published work.
2 Related Work
Lack of diversity in text may result from repetition of lengthy strings or owe to subtle distributional patterns (Holtzman et al., 2019; Meister et al., 2022, 2023a). We focus on scores that aim to capture overt repetition across outputs, and leave for future work similar analysis of semantic and structural diversity scores (Bär et al., 2012; Shaib et al., 2024).
Conditional generation tasks such as image captioning and dialog summarization have offered observations regarding the diversity of produced texts. Prior work on both of these tasks has documented that models tend to repeat the same text for different contexts. Li et al. (2016) find that four phrases account for about a third of all turns produced by a conversational agent, and Devlin et al. (2015) report that more than half of automatically generated captions are repeated verbatim for different images. Self-repetition (Salkar et al., 2022), i.e., exact repetition of the same $n$-gram ($n\geq 4$) across different summaries is a practical way of measuring repetition for tasks where generated text is lengthier. In such cases exact matches rarely occur but repetition is common, especially relative to the training data (Wang et al., 2023a).
Above we discussed a variety of metrics but it is unclear which of these to use when, and how to efficiently visualize the implications of lower diversity in tasks that comprise only text. Further, prior work has shown that human judgments of diversity are difficult to reliably collect. Humans tend to implicitly conflate quality of text with its diversity, and it can be difficult to separate content and lexical diversity in these assessments (Tevet and Berant, 2021). Motivated by the lack of visualization platforms and the need to provide better methods for human evaluations of repetition in large datasets, we build an interactive tool that allows users to browse highlighted instances of “lower diversity” text, as shown in Figure 1.
2.1 A Smorgasbord of Text Diversity Scores
@@
Scores used to measure diversity across a corpus of texts derive from two core ideas: Computing average similarity between pairs of outputs produced by the same model for different inputs, and computing variants of token/type ratio. The former are adapted from common approaches to reference-based text generation using standard measures of pairwise similarity; the latter track the diversity of vocabulary measured as the ratio of unique words to total words produced, with outputs from a model concatenated into a single text. We first describe each score, and then present insights regarding their mutual redundancy. We also consider run-times, which are lengthy for some metrics and may render them impractical for analysis of a large datasets. All scores are defined for a set of generated texts $D$, each conditioned on its respective input.
@@
The quality of text in machine translation, summarization, and image captioning is often reported in terms of overlap with a reference text. This idea can be adapted to measure diversity across different outputs by using one generated text as a “reference” and measuring the similarity of other outputs against this. Self-BLEU measures similarity between all text pairs in $D$ using BLEU (Zhu et al., 2018). BLEU could be replaced with other similarity scores, e.g., ROUGE-L or BERTScore. These variants are called homogenization scores and have recently been used to compare the diversity of texts produced under several conditions (Padmakumar and He, 2023a).
@@
All homogenization scores calculate an aggregate similarity across pairs of examples (Equation 1). Here the similarity score of choice is ROUGE-L Lin and Och 2004, which quantifies overlap in terms of longest common sub-sequences between all pairs of text in a corpus instead of the fixed $n$-gram size used in other ROUGE variants:
@@
Self-repetition measures the tendency of LMs to repeat long $n$-grams across different outputs (Salkar et al., 2022).
@@
The token-type ratio for a text is the unique token count divided by the total count of tokens. Moving Average Token Type Ratios (MATTRs) measures the lexical dynamics across a text which is insensitive to text length. This captures the repetition of a given word in segments of text and does not explicitly account for longer repeated sequences (Covington and McFall, 2010).
@@
NGD extends the idea of token-type ratio to longer $n$-grams (Padmakumar and He, 2023a; Meister et al., 2023b; Li et al., 2023). It is defined as a ratio of unique $n$-gram counts to all $n$-gram counts:
@@
The probability of text under a Hypergeometric Distribution D (HD-D) is an another measure of lexical diversity (McCarthy and Jarvis, 2010). [^5] For both HD-D and MATTR, we use the implementation provided in the lexical-diversity package (https://pypi.org/project/lexical-diversity/). HD-D does not capture repetition of sub-sequences.
2.2 Compression Ratios for Diversity
@@
The diversity scores introduced so far are all a function of the number of repeated substrings across outputs. We use gZip to compress the concatenated text of all outputs generated by a model. The compression ratio is then the ratio between the size of the compressed file to that of the original file. High compression ratios imply more redundancy:
@@
3 Evaluating Repetition with diversity
3.1 Design of the Diversity Package
The diversity package uses NLTK (Bird and Loper, 2004) and SpaCY Honnibal et al. (2020) to tag text with parts of speech and extract $n$-grams. Users can install the package via pip (assuming Python 3.10+). They can then calculate various diversity metrics over a list of texts as follows:
1from diversity import compression_ratio,
2 homogenization_score,
3 ngram_diversity_score,
4 get_pos
5
6text = [
7"I enjoy walking with my cute dog for the rest of the day, ...",
8"I enjoy walking with my cute dog. The only time I felt ...",
9"I enjoy walking with my cute dog, and I really enjoy running..."
10]
11
12# compression ratios
13cr = compression_ratio(text, ’gzip’)
14cr_pos = compression_ratio(get_pos(text)[1], ’gzip’)
15
16# homogenization scores
17hs_rougel = homogenization_score(text, ’rougel’)
18hs_bert = homogenization_score(text, ’bertscore’)
19self_bleu = homogenization_score(text, ’bleu’)
20
21# other
22self_rep = self_repetition_score(text)
23nds = ngram_diversity_score(text, n=4)
Users can also extract repeated text based on part of speech ngrams with the extract_patterns and match_patterns functions.
1n = 5
2top_n = 100
3extract_patterns(text, n, top_n)
First, the user extracts patterns using the extract_patterns function by specifying the n-gram length to search for, and the top_n most repeated n-grams to return. This returns a Python dictionary where the keys are the part-of-speech n-grams and the values are the raw text n-grams matching those patterns. The pattern matches are based on the frequency seen across the entire dataset, i.e., a part-of-speech pattern is only a pattern if it appears in more than 2 texts in the original input. Default values consider the top 100 part-of-speech patterns (sorted by frequency).
Then, using match_patterns, a user can identify all patterns in a single text from the input:
1idx = 2
2match_patterns(text[idx], patterns)
which returns a list of tuples containing the pattern in the first position, and the matched substring in the second position.
Many diversity metrics require pairwise comparisons. With larger datasets, this can become infeasible to compute (see Appendix B). We implement a few methods to increase efficiency: We cache to store already computed pairs, and on the website we chunk incoming data and compute pattern searches over batches.
3.2 Design of the Web UI
The diversity Web UI functionality is built around the Python package. All front-end design uses Vue.js and Tailwind CSS.
The UI offers the same functionality as the package, however allows users to measure and explore repetition without requiring any Python knowledge. Figure 2 shows the main pages of the site: users can begin by (A) either uploading their own text file for analysis or selecting one of the demo datasets provided on the site. Then, the user is prompted to select one of three types of analyses: either (B) to explore part-of-speech patterns, (C) to explore verbatim repeated text, or (D) to measure various diversity metrics of the dataset. Datasets are processed upon upload, and nothing is stored on the backend server aside from the existing demo datasets.
(B) Templates
The templates tab allows users to explore extracted part-of-speech $n$-grams in their selected dataset. The left-most column displays pattern length of $n=[2,10]$. The user can then scroll through all of the templates, select some or all, and see the highlighted text in the middle panel corresponding to the template. The templates are assigned a colour when selected to indicate the corresponding matched text. The right-most column provides a reference for all the part-of-speech tags from SpaCY.  [^7] https://spacy.io/usage/linguistic-features The default pattern length is set to $n=4$. Other lengths will load when selected.
(C) Exact Match
The exact matches tab allows a user to explore exact text matches in their dataset. The top provides two sliders: the left slider allows the user to set a string length to search for ($n=[2,10]$), and the right the minimum number of documents in which the string must appear ($n=[2,10]$). The minimum document occurrence slider defaults to 2. Once selected, the user can scroll through to see the repeated text in bold, and the full document text in which the string appears, as well as the number of documents.
(D) Diversity Metrics
The diversity metrics tabs reports the recommended metrics from our evaluation: Compression Ratio, POS Compression, Self-BLEU, Self-Repetition, and Homogenization with BERTScore. We display these values alongisde a guide to the metrics on the right-hand side.
3.3 Use-Cases
Our implementation of compression ratios over parts-of-speech and tokens (along with BERTScore, Self-BLEU, and self-repetition) have been used in prior works (by other groups) to evaluate diversity in model evaluation and alignment (Lake et al., 2024; Moon et al., 2024; Fernandez et al., 2024), and for reporting diversity over synthetic datasets (Chang et al., 2024; Hastings et al., 2024). Due to its computational efficiency, compression ratios have also been used as optimization parameters in decoding strategies (Lanchantin et al., 2025).
Shaib et al. (2024) use the pattern analysis in diversity to measure and evaluate the prevalence of syntactic patterns in LLMs. Similarly, Wadhwa et al. (2025) extract part-of-speech patterns in distillation tasks for model attribution. Further, insights from our evaluation of diversity metrics have been used to inform how to report diversity with respect to text length and data sizes (Guo et al., 2023; Hastings et al., 2024).
Overall, prior work has shown how NLP researchers can greatly benefit from open access to diversity as well as the evaluation conducted across the implemented metrics. By releasing a user interface, we broaden use-cases for users who would like to explore patterns and diversity in text datasets without needing to use Python.
4 Platform Evaluation: Comparative Analysis of Diversity Metrics
4.1 Data and Models
@@
We compute diversity scores for the outputs of six instruction tuned models on the CNN/DailyMail (Hermann et al., 2015) and XSUM (Narayan et al., 2018) English news summarization datasets: Llama-2 (Touvron et al., 2023a), GPT-4 (OpenAI, 2023), FlanT5-XXL (Longpre et al., 2023), StableLM (Taori et al., 2023; Chiang et al., 2023; Anand et al., 2023), Mistral (Jiang et al., 2023), and StableBeluga (Touvron et al., 2023b; Mukherjee et al., 2023). [^8] All models—except GPT-4—downloaded from HuggingFace (https://huggingface.co/models). We selected these models to cover a range of availability (open and closed), and architectures (encoder-decoder, decoder-only). [^9] We use prompts for summarization provided by each model, where available. See Appendix C.4. The lengths of texts vary considerably by source, for reference and model-produced text alike, so we also note average lengths when reporting diversity.
@@
5 Text Length as a Confounder
To keep compute time and costs manageable, we randomly sample 500 inputs from CNN/DailyMail and XSUM for analysis. Table 2 reports diversity scores for outputs generated by the six LLMs for these inputs. Table 2 (top) reports scores for human-written texts: The article given as input for summarization, the baseline summary comprising the first three sentences of the news article, and the reference summary. These scores serve as a reference point for the diversity scores of the models.